# Stability-Aware Training of Neural Network Interatomic Potentials with Differentiable Boltzmann Estimators

### v1

@@ A.1 Derivation of Boltzmann Estimator
$\mathcal{E}(\Gamma^{(1)},\ldots,\Gamma^{(N)})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\Gamma)\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\right].$
We note that a similar estimator is obtained in [undefaal] by differentiating through a reweighting scheme arising from thermodynamic perturbation theory. We have presented an alternative derivation that does not require reweighting.
A.2 Localized Boltzmann Estimator
We provide additional details on the Localized Boltzmann Estimator. Formally, define a neighborhood of $n$ atoms $\mathcal{N}=\{x_{1},x_{2},\ldots,x_{n}\mid x_{i}\in\mathbb{Z},1\leq x_{i}\leq N\ \text{for all}\ i=1,2,\ldots,n\}$. The local state $\gamma$ is defined as $\gamma=\{[r_{\mathcal{N}_{1}};\ldots;r_{\mathcal{N}_{n}}],[P_{\mathcal{N}_{1}};\ldots;P_{\mathcal{N}_{n}}]\}$. The local energy is then defined as $U_{\theta}(\gamma)=\sum_{i=1}^{n}U_{\theta}(\gamma_{i})$, where $\gamma_{i}=\{r_{\mathcal{N}_{i}},P_{\mathcal{N}_{i}}\}$ contains the position and momenta of the $i^{th}$ atom in the local neighborhood. We can obtain this by noting that NNIPs first compute individual atomic energies before summing over all atoms to obtain the global energy. Finally, given $N$ i.i.d. samples of local states $\gamma^{(1)},\ldots,\gamma^{(N)}$, where each $\gamma^{(i)}$ is extracted from a global state $\Gamma^{(i)}\sim P_{\theta}(\Gamma)$, we define the $N$-sample localized estimator of the Jacobian $\frac{\partial\mathbb{E}_{\gamma\sim P_{\theta}(\gamma)}[g(\gamma)]}{\partial\theta}$ as
$\mathcal{E}(\gamma^{(1)},\ldots,\gamma^{(N)})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\gamma)\cdot\nabla_{\theta}U_{\theta}(\gamma)}^{\top}\right]\right]$ (8)
The localized estimator follows from the original Boltzmann estimator due to the fact that as a subset of the larger Boltzmann-distributed system, any local neighborhood also obeys a Boltzmann-distribution. In practice, we extract multiple local neighborhoods from each global state to increase the sample size and state space coverage. See Supplementary Section A.6 for a description of the neighborhood extraction procedure used for the Water system.
@@ A.3 StABlE Training Algorithm
We provide an algorithmic description of our StABlE-Training procedure. The algorithm proceeds in two main phases: simulation, which seeks out unstable regions of molecular phase space, and learning, which corrects the instability via supervision from the reference observables.
@@
[…]
@@
15:   $\bar{\Gamma}_{curr}\leftarrow\{\Gamma^{(1)}({T_{f_{1}}}),\Gamma^{(2)}({T_{f_{2}}}),\ldots\Gamma^{(R)}({T_{f_{R}}})\}$
@@
19:  // $\bar{\Gamma}_{curr}$ now contains $\{\Gamma^{(1)}({T_{f_{1}}}),\Gamma^{(2)}({T_{f_{2}}}),\ldots\Gamma^{(R)}({T_{R}})\}$, where $T_{f_{i}}$’s are per-replica total simulation times
20:  // At least $f_{max}$ fraction of replicas are unstable at this point Learning Phase
21:  while $f_{unst}>f_{min}$ do
22:   Rewind all trajectories by $t$ timesteps: $\bar{\Gamma}_{curr}\leftarrow\{\Gamma^{(1)}({T_{f_{1}}-t}),\ldots,\Gamma^{(R)}({T_{f_{R}}-t})\}$
23:   Simulate all replicas w/ $U_{\theta}$ for $t$ steps starting from $\bar{\Gamma}_{curr}$
24:   Update $f_{unst}$
25:   Compute observables $\{\widehat{\mathbb{E}}_{\Gamma}[g^{(i)}(\Gamma)]\}_{i=1}^{N}$ over all length-$t$ trajectories
26:   $\mathcal{L}_{obs}\leftarrow\sum_{i=1}^{N}\|\widehat{\mathbb{E}}_{\Gamma}[g^{(i)}(\Gamma)]-g_{\text{ref}}^{(i)}\|^{2}$
27:   $\mathcal{L}_{\text{QM}}\leftarrow$ energy and forces loss of $U_{\theta}$ on dataset $\mathcal{D}_{\text{train}}$
28:   $\theta\leftarrow\theta-\alpha\cdot\nabla_{\theta}(\mathcal{L}_{obs}+\lambda\mathcal{L}_{\text{QM}})$ // Compute Boltzmann Estimator
29:  end while
30:  Mark stable replicas active, unstable replicas inactive for next simulation phase
31:  // At most $f_{min}$ fraction of replicas are unstable at this point
32:until Convergence or maximum cycles reached
@@ A.4 Observables
where $\delta$ is the Dirac-Delta function. Although the observable need not be differentiable with respect to $r$ for our Boltzmann learning framework, we compute a differentiable version of $h(r)$ via Gaussian smearing as in [undefaar] to facilitate comparison with differentiable simulation methods.
Radial Distribution Function.
@@
where $t_{0}$ is an initial time, $v_{i}(t)$ is the velocity of the $i^{th}$ atom at timestep $t$, $\cdot$$\cdot$> is an inner product, and $S$ is the total number of samples considered given the summations over initial times and atoms. In this work, we compute the VACF over a window of 100 consecutive simulation timesteps, and normalize the values by the autocorrelation at $\Delta t=0$ to restrict the range to $[0,1]$.
Diffusivity Coefficient
@@ A.5 Stability Criteria
Maximum Bond Length Deviation.
Adapted from [undefv], the maximum bond length deviation metric captures unphysical bond stretching or collapse in small flexible molecules. According to this criterion, a simulation becomes unstable at time $T$ if,
@@
where $\mathcal{B}$ is the set of all bonds, $i,j$ are the two endpoint atoms of the bond, and $b_{i,j}$ is the equilibrium bond length computed from the reference simulation. Following [undefv], we set $\Delta=0.5A$ for final stability evaluation. However, we adopt a more conservative value of $\Delta=0.25A$ during training in order to detect and correct instability earlier. We use this criterion for the MD17 and MD22 datasets.
Minimum Intermolecular Distance.
@@
Radial Distribution Function Mean Absolute Error.
The minimum intermolecular distance metric is appropriate at train-time to detect instability early before it cascades to the rest of the system. However, it is too sensitive to use for evaluation, as realistic simulation can still be achieved for some time after the occurrence of a highly localized instability. Therefore, following [undefv], we adopt an instability metric based on the radial distribution function, defined as,
$\int_{r=0}^{\infty}\left\|RDF_{\text{ref}}(r)-\left\langle RDF^{t}(r)\right\rangle_{t=T}^{T+\tau}\right\|\,dr>\Delta,$
where $\langle\cdot\rangle$ is the averaging operator, $\tau$ is a short time window, and $\Delta$ is the stability threshold. We use $\tau=1$ ps and $\Delta=3.0$ for water. The stability criterion is triggered if any of the three element-conditioned water RDFs (H-H, O-O, or H-O) exceeds the threshold.
Justification of Chosen Stability Thresholds.
Following [undefv], we choose stability thresholds such that a realistic, high-fidelity simulation at the chosen temperature would virtually never cross the threshold. This means that if a simulation does cross the threshold, this is indicative of catastrophic failure. Thresholds are set more conservatively during training in order to facilitate early detection of potential collapse. Figure 7 shows the distribution of values of the stability criterion over high-fidelity reference simulations for the three systems considered in this work, along with thresholds chosen to denote instability for training and evaluation.
@@ A.6 Architecture and Training Details
We provide details on the model architectures and training procedures used in this work. MD simulations and NNIPs are written in the PyTorch framework and are built upon the MDsim [undefv] and Atomic Simulation Environment [undefam] packages. All training is performed on a single NVIDIA A100 GPU.
Table 1 provides details on the NNIP architectures. $r_{max}$ is the cutoff distance used to construct the radius graph. $l_{max}$ denotes the level of E(3) equivariance used in the network.
@@
SchNet [undefaae] E(3)-invariant 0.12M 5.0 -
NequIP [undefd] E(3)-equivariant 0.12M 5.0 1
GemNet-T [undefw] SE(3)-equivariant 1.89M 5.0 -
Table 1: NNIP Architecture Details.
For energy and forces pre-training, we follow the protocols in [undefv]. In order to isolate the effect of observable-based learning, we begin StABlE Training only after pre-training has fully converged (that is, when $\mathcal{L}_{\text{QM}}$ has reached a plateau). This means that any improvements in stability or accuracy as a result of StABlE Training can be attributed to the learning signal from the reference observable, as opposed to the regularization from the QM energy and forces data.
We include relevant settings used for StABlE Training in Table 2. $\alpha$ is the learning rate, $\lambda$ is the strength of energy and forces regularization, $t$ is the number of simulation timesteps per epoch, and $R$ is the number of parallel replicas. We note that in practice, we compute the outer products and empirical means in the Boltzmann estimator in batched fashion. Thus, to limit memory usage, we compute $\frac{N}{B}$ separate Boltzmann estimators from minibatches of $B<N$ states and subsequently average them to produce a final estimator.
NNIP Stability Training Estimator $\alpha$ $\lambda$ t $R$ $B$
@@
Details on Local Neighborhood Sampling Scheme for Water.
For water training, from each sampled global state we consider all possible local neighborhoods containing a single water molecule. Since instability is highly localized, the majority of local neighborhoods are stable, leading to an imbalance in the sampled states. To provide more informative samples for optimization, we assign the individual sampled molecules to one of 100 bins based on their deviation from the equilibrium O-H bond length. Then, we choose a fixed number (in this case 5) molecules from each bin from which to compute the final localized estimator.
@@
In Table 3, we provide the total wall clock time spent on QM pre-training, as well as the subsequent StABlE Training. All runtimes were measured on an NVIDIA A100 GPU. We note that especially for Ac-Ala3-NHMe and Water, StABlE Training incurs a relatively small marginal computational cost beyond that of QM pre-training.
@@ A.7 Simulation Details
We provide MD simulation details in Table 4. All systems considered in this work are simulated with a Nose-Hoover thermostat. All simulation parameters are chosen to be consistent with prior literature [undefq, undefp, undefv].
Temperature Nose-Hoover Timestep Periodic Boundary
(K) Effective Mass (cm-1) (fs) Conditions
Aspirin 500 2000 0.5 No
Ac-Ala3-NHMe 500 2000 0.5 No
Water 300 2000 1 Yes
@@ A.8 Evaluation Details
We provide further details on the protocol used to evaluate the NNIPs considered in this work. Our evaluation protocol is centered around MD simulations. To facilitate direct comparison between StaBlE-trained NNIPs and those trained only on energy and forces reference data, for a given molecular system we perform MD simulations starting from the same initial configurations for all models. We choose the number of parallel replicas and total simulation time on a per-system basis so as to saturate GPU memory usage while remaining within a reasonable computational budget. Simulation conditions are the same as described in Section A.7 except for the temperature generalization experiment (Section 3.2), in which the temperature of simulation is varied. Table 5 summarizes the relevant evaluation parameters for each system.
@@
Aspirin 256 300 Bond Length Deviation, 0.5 A
@@ A.9 Temperature-Reweighting of Observables
We provide further details on the reweighting process used to estimate the reference distribution of interatomic distances at 350K and 700K.
Under the canonical ensemble, microstates follow a Boltzmann distribution $P_{\theta}(\Gamma)\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}{C(\theta)}$. Consider states $\Gamma^{(1)},\ldots,\Gamma^{(N)}$ sampled at temperature $T_{1}$. Define a reweighting factor for each sample as follows,
$w_{i}$ $=\frac{\frac{P_{\theta}(\Gamma;T_{1})}{P_{\theta}(\Gamma;T_{2})}}{\sum_{i=1}^{N}\frac{P_{\theta}(\Gamma^{i};T_{1})}{P_{\theta}(\Gamma^{i};T_{2})}}$ $=\frac{\exp(-\frac{\mathcal{H}_{\theta}(\Gamma^{i})}{k_{B}}(\frac{1}{T_{2}}-\frac{1}{T_{1}}))}{\sum_{i=1}^{N}\exp(-\frac{\mathcal{H}_{\theta}(\Gamma^{i})}{k_{B}}(\frac{1}{T_{2}}-\frac{1}{T_{1}}))}.$
We can then compute a reweighted Monte Carlo estimate of the observable at $T_{2}$ as follows [undefaal],
$g_{true,T_{2}}=\sum_{i=1}^{N}w_{i}g(\Gamma^{(i)})$
The statistical error of the reweighted Monte Carlo estimate is captured by the effective sample size, $N_{eff}\approx e^{-\sum_{i=1}^{N}w_{i}ln(w_{i})}$ [undefl]. A small effective sample size indicates that a few samples with high weights dominate the average; this occurs for large differences between $T_{1}$ and $T_{2}$. To select lower and upper temperatures at which to perform the temperature generalization experiment, we set a minimum $N_{eff}=1000$, leading us to choose 350K and 700K (Supplemental Figure 8).
@@
A.10 Analysis of Energy and Forces Errors
We study the effect of two hyperparameters of StABlE Training, the learning rate $\alpha$ and the strength of QM regularization $\lambda$, on the energy and forces errors of a SchNet NNIP on a held-out test set of aspirin structures. We perform StABlE Training as described in Section 2 for learning rates ranging from $10^{-5}$ to $10^{-3}$ and QM regularization coefficients ranging from $10^{0}$ to $10^{2}$. We perform evaluation of each trained model via MD simulation of 256 parallel replicas at 500K.

### v2

@@ A.1 Derivation of Boltzmann Estimator
$\mathcal{E}(\Gamma_{1},\ldots,\Gamma_{N})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\Gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\Gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\Gamma)\nabla_{\theta}U_{\theta}(\Gamma)}^{\top}\right]\right].$
We note that a similar estimator is obtained in Thaler & Zavadlav (2021) by differentiating through a reweighting scheme arising from thermodynamic perturbation theory. We have presented an alternative derivation that does not require reweighting.
A.2 Extension to Other Statistical Ensembles
The Boltzmann Estimator is applicable out-of-the-box to any statistical ensemble where the probability of a microstate can be written as $P_{\theta}(\Gamma)\propto\exp\left(-\frac{1}{k_{B}T}\left[\mathcal{H}_{\theta}(\Gamma)+\mathcal{X}(\Gamma)\right]\right)$, where $\mathcal{X}$ contains state-dependent thermodynamic variables. For the isothermal-isobaric (NPT) ensemble, $\mathcal{X}(\Gamma)=pV(\Gamma)$, where $p$ is the simulation pressure and $V(\Gamma)$ is the volume of the microstate. For the grand canonical ($\mu VT$) ensemble, $\mathcal{X}(\Gamma)=\mu N(\Gamma)$, where $\mu$ is the chemical potential and $N(\Gamma)$ is number of particles in the microstate. For the canonical (NVT) ensemble considered in this work, $\mathcal{X}(\Gamma)=0$. Since $\mathcal{X}(\Gamma)$ is independent of the MLFF parameters $\theta$ in all cases, it can effectively be absorbed into the kinetic energy component of the Hamiltonian, which does not affect the computation of the estimator. Our derivation thus proceeds in the same fashion and yields the same estimator. The Localized Boltzmann Estimator also holds as before.
@@ A.3 StABlE Training Algorithm
We provide an algorithmic description of our StABlE-Training procedure.
@@
7:$\bar{\Gamma}_{curr}\leftarrow\{\Gamma_{1}(0),\Gamma_{2}(0),\ldots\Gamma_{R}(0)\}\sim\mathcal{D}_{\text{train}}$
@@
15:         $\bar{\Gamma}_{curr}\leftarrow\{\Gamma_{1}({T_{f_{1}}}),\Gamma_{2}({T_{f_{2}}}),\ldots\Gamma_{R}({T_{f_{R}}})\}$
@@
19:     // $\bar{\Gamma}_{curr}$ now contains $\{\Gamma_{1}({T_{f_{1}}}),\Gamma_{2}({T_{f_{2}}}),\ldots\Gamma_{R}({T_{R}})\}$, where $T_{f_{i}}$’s are per-replica total simulation times. At least $f_{max}$ fraction of replicas are unstable at this point Learning Phase
20:     while $f_{unst}>f_{min}$ do
21:         Rewind all trajectories by $t$ timesteps: $\bar{\Gamma}_{curr}\leftarrow\{\Gamma_{1}({T_{f_{1}}-t}),\ldots,\Gamma_{R}({T_{f_{R}}-t})\}$
22:         Simulate all replicas w/ $U_{\theta}$ for $t$ steps starting from $\bar{\Gamma}_{curr}$
23:         Update $f_{unst}$
24:         Compute observables $\{\widehat{\mathbb{E}}_{\Gamma}[g^{(i)}(\Gamma)]\}_{i=1}^{N}$ over all length-$t$ trajectories
25:         $\mathcal{L}_{obs}\leftarrow\sum_{i=1}^{N}\|\widehat{\mathbb{E}}_{\Gamma}[g^{(i)}(\Gamma)]-g_{\text{ref}}^{(i)}\|^{2}$
26:         $\mathcal{L}_{\text{QM}}\leftarrow$ energy and forces loss of $U_{\theta}$ on dataset $\mathcal{D}_{\text{train}}$
27:         $\theta\leftarrow\theta-\alpha\cdot\nabla_{\theta}(\mathcal{L}_{obs}+\lambda\mathcal{L}_{\text{QM}})$ // Compute Boltzmann Estimator
28:     end while
29:     Mark stable replicas active, unstable replicas inactive for next simulation phase
30:     // At most $f_{min}$ fraction of replicas are unstable at this point
31:until Convergence or maximum cycles reached
@@ A.4 Observables
where $\delta$ is the Dirac-Delta function. Although the observable need not be differentiable with respect to $r$ for our Boltzmann learning framework, we compute a differentiable version of $h(r)$ via Gaussian smearing as in Wang et al. (2023a) to facilitate comparison with differentiable simulation methods.
@@
where $t_{0}$ is an initial time, $v_{i}(t)$ is the velocity of the $i^{th}$ atom at timestep $t$, <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > is an inner product, and $S$ is the total number of samples considered given the summations over initial times and atoms. In this work, we compute the VACF over a window of 100 consecutive simulation timesteps, and normalize the values by the autocorrelation at $\Delta t=0$ to restrict the range to $[-1,1]$.
@@ A.5 Stability Criteria
Adapted from Fu et al. (2022), the maximum bond length deviation metric captures unphysical bond stretching or collapse in small flexible molecules. According to this criterion, a simulation becomes unstable at time $T$ if,
@@
where $\mathcal{B}$ is the set of all bonds, $i,j$ are the two endpoint atoms of the bond, and $b_{i,j}$ is the equilibrium bond length computed from the reference simulation. Following Fu et al. (2022), we set $\Delta=0.5A$ for final stability evaluation. However, we adopt a more conservative value of $\Delta=0.25A$ during training in order to detect and correct instability earlier. We use this criterion for the MD17 and MD22 datasets.
@@
The minimum intermolecular distance metric is appropriate at train-time to detect instability early before it cascades to the rest of the system. However, it is too sensitive to use for evaluation, as realistic simulation can still be achieved for some time after the occurrence of a highly localized instability. Therefore, following Fu et al. (2022), we adopt an instability metric based on the radial distribution function, defined as,
[…]
@@ A.6 Architecture and Training Details
We provide details on the model architectures and training procedures used in this work. MD simulations and MLFFs are written in the PyTorch framework and are built upon the MDsim Fu et al. (2022) and Atomic Simulation Environment Larsen et al. (2017) packages. All training is performed on a single NVIDIA A100 GPU.
Supplementary Table 1 provides details on the MLFF architectures. $r_{max}$ is the cutoff distance used to construct the radius graph. $l_{max}$ denotes the level of E(3) equivariance used in the network.
@@
SchNet Schütt et al. (2018) E(3)-invariant 0.12M 5.0 -
NequIP Batzner et al. (2022) E(3)-equivariant 0.12M 5.0 1
GemNet-T Gasteiger et al. (2021) SE(3)-equivariant 1.89M 5.0 -
Table 1: MLFF Architecture Details.
For energy and forces pre-training, we follow the protocols in Fu et al. (2022). In order to isolate the effect of observable-based learning, we begin StABlE Training only after pre-training has fully converged (that is, when $\mathcal{L}_{\text{QM}}$ has reached a plateau). This means that any improvements in stability or accuracy as a result of StABlE Training can be attributed to the learning signal from the reference observable, as opposed to the regularization from the QM energy and forces data.
We include relevant settings used for StABlE Training in Supplementary Table 2. $\alpha$ is the learning rate, $\lambda$ is the strength of energy and forces regularization, $t$ is the number of simulation timesteps per epoch, and $R$ is the number of parallel replicas. We note that in practice, we compute the outer products and empirical means in the Boltzmann estimator in batched fashion. Thus, to limit memory usage, we compute $\frac{N}{B}$ separate Boltzmann estimators from minibatches of $B<N$ states and subsequently average them to produce a final estimator.
MLFF Stability Training Estimator $\alpha$ $\lambda$ t $R$ $B$
@@
In Supplementary Table 3, we provide the total wall clock time spent on QM pre-training, as well as the subsequent StABlE Training. All runtimes were measured on an NVIDIA A100 GPU. We note that especially for Ac-Ala3-NHMe and Water, StABlE Training incurs a relatively small marginal computational cost beyond that of QM pre-training.
@@ A.7 Simulation Details
We provide MD simulation details in Supplementary Table 4. During training, all systems are simulated with a Nose-Hoover thermostat. During evaluation, either a Nose-Hoover or Langevin thermostat is used based on whichever one yields better stability. Thermostat parameters are chosen to be consistent with prior literature Chmiela et al. (2017; 2022); Fu et al. (2022): for Nose-Hoover simulations, the temperature coupling constant is set to 20 fs, and for Langevin simulations, the friction coefficient is set to 0.1 ps-1.
Temperature Timestep Periodic Boundary Simulation
(K) (fs) Conditions Thermostat
Aspirin 500 0.5 No Langevin
Ac-Ala3-NHMe 500 0.5 No Nose-Hoover
Water 300 1 Yes Nose-Hoover
@@
For simulating water in the NPT ensemble at 300K and 1 atm, we employ the Berendsen barostat with temperature and pressure coupling times of 20 fs and 2 ps respectively. We limit the per-timestep change in momentum and volume to 10% and 1% respectively to prevent instabilities resulting from large fluctuations. We found the Berendsen barostat to be more stable than combined Parinello-Rahman and Nose-Hoover dynamics, which would be the more conventional choice for NPT simulations with standard potentials. We speculate that this may be due to qualitative differences in the behavior of ML potentials compared to classical or ab-initio potentials, and these differences warrant further investigation in the future.
@@ A.8 Evaluation Details
We provide further details on the protocol used to evaluate the MLFFs considered in this work. Our evaluation protocol is centered around MD simulations. To facilitate direct comparison between StABlE-trained MLFFs and those trained only on energy and forces reference data, for a given molecular system we perform MD simulations starting from the same initial configurations for all models. We choose the number of parallel replicas and total simulation time on a per-system basis so as to saturate GPU memory usage while remaining within a reasonable computational budget. Simulation conditions are the same as described in Section A.7 except for the temperature generalization experiment, in which the temperature of simulation is varied. Supplementary Table 5 summarizes the relevant evaluation parameters for each system.
@@
Aspirin 256 1000 Bond Length Deviation, 0.5 A
@@
Justification of Chosen Stability Thresholds.
Following Fu et al. (2022), we choose stability thresholds such that a realistic, high-fidelity simulation at the chosen temperature would virtually never cross the threshold. This means that if a simulation does cross the threshold, this is indicative of catastrophic failure. Thresholds are set more conservatively during training in order to facilitate early detection of potential collapse. Supplementary Figure 7 shows the distribution of values of the stability criterion over high-fidelity reference simulations for the three systems considered in this work, along with thresholds chosen to denote instability for training and evaluation.
@@ A.9 Temperature-Reweighting of Observables
We provide further details on the reweighting process used to estimate the reference distribution of interatomic distances at 350K and 700K. This was used in the temperature generalization experiments described in Section 4.2.
Under the canonical ensemble, microstates follow a Boltzmann distribution $P_{\theta}(\Gamma)\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{\exp\left(-\frac{1}{k_{B}T}\mathcal{H}_{\theta}(\Gamma)\right)}{C(\theta)}$. Consider states $\Gamma_{1},\ldots,\Gamma_{N}$ sampled at temperature $T_{1}$. Define a reweighting factor for each sample as follows,
$w_{i}$ $=\frac{\frac{P_{\theta}(\Gamma;T_{1})}{P_{\theta}(\Gamma;T_{2})}}{\sum_{i=1}^{N}\frac{P_{\theta}(\Gamma_{i};T_{1})}{P_{\theta}(\Gamma_{i};T_{2})}}$ $=\frac{\exp(-\frac{\mathcal{H}_{\theta}(\Gamma_{i})}{k_{B}}(\frac{1}{T_{2}}-\frac{1}{T_{1}}))}{\sum_{i=1}^{N}\exp(-\frac{\mathcal{H}_{\theta}(\Gamma_{i})}{k_{B}}(\frac{1}{T_{2}}-\frac{1}{T_{1}}))}.$
We can then compute a reweighted Monte Carlo estimate of the observable at $T_{2}$ as follows Thaler & Zavadlav (2021).
$g_{true,T_{2}}=\sum_{i=1}^{N}w_{i}g(\Gamma_{(i)})$
The statistical error of the reweighted Monte Carlo estimate is captured by the effective sample size, $N_{eff}\approx e^{-\sum_{i=1}^{N}w_{i}ln(w_{i})}$ Carmichael & Shell (2012). A small effective sample size indicates that a few samples with high weights dominate the average; this occurs for large differences between $T_{1}$ and $T_{2}$. To select lower and upper temperatures at which to perform the temperature generalization experiment, we set a minimum $N_{eff}=1000$, leading us to choose 350K and 700K (Figure 8).
A.10 StABlE Training in the Isothermal-Isobaric Ensemble
We repeat StABlE Training on the all-atom water system, and this time simulate in the isothermal-isobaric (NPT) ensemble with a temperature of 300K and a pressure of 1 atm. We use the same training settings outlined in Supplementary Table 2. We find similar results to when we simulate in the canonical (NVT) ensemble: StABlE Training yields clear stability improvements, increasing the median stable simulation time from 51 to 165 picoseconds. The stability and quality of estimated observables is slightly lower than in NVT simulations, including some unphysical collisions in the short-range region of the element-conditioned RDFs. This may be due to the distribution shift induced by the continuously changing box size in NPT simulations, which was not seen during pretraining of the GemNet-T potential.
Figure 9: Results of StABlE Training a GemNet-T model for all-atom water simulation in the isothermal-isobaric (NPT) ensemble. StABlE Training yields considerable improvements in stable simulation time, and maintains or slightly improves the accuracy of recovered observables.
A.11 Analysis of Energy and Forces Errors
We study the effect of two hyperparameters of StABlE Training, the learning rate $\alpha$ and the strength of QM regularization $\lambda$, on the energy and forces errors of a SchNet MLFF on a held-out test set of aspirin structures. We perform StABlE Training for learning rates ranging from $10^{-5}$ to $10^{-3}$ and QM regularization coefficients ranging from $10^{0}$ to $10^{2}$. We perform evaluation of each trained model via MD simulation of 256 parallel replicas at 500K.